OpenStudy (anonymous):

For the standard form of a circle (x –9)^2+(y+3)^2=64, a.Find the center:(h,k)= ? b.Find the radius:r= ?

OpenStudy (anonymous):

Center is (h,k) in the equation of a circle (x-h)^2 + (y-k)^2 = r^2 so the center will be at (9,-3)

OpenStudy (anonymous):

r^2 = 64 so r, the radius = 8